# Projective Cohen-Macaulay monomial curves and their affine charts

### v2

@@ Introduction
Now consider an integer $d>0$ and a sequence $a_{0}=0<a_{1}<\cdots<a_{n}=d$ of relatively prime integers. Denote by $\mathcal{C}$ the projective monomial curve $\mathcal{C}\subset\mathbb{P}_{k}^{\,n}$ of degree $d$ parametrically defined by $x_{i}=u^{a_{i}}v^{d-a_{i}}$ for all $i\in\{0,\ldots,n\}$, i.e., $\mathcal{C}$ is the Zariski closure of
@@
Taking $\mathcal{A}=\{\mathbf{a}_{0},\ldots,\mathbf{a}_{n}\}\subset\mathbb{N}^{2}$ with $\mathbf{a}_{i}=(a_{i},d-a_{i})$ for all $i=0,\ldots,n$, one has that $I_{\mathcal{A}}$ is the vanishing ideal of $\mathcal{C}$, and the coordinate ring of $\mathcal{C}$ is the two-dimensional ring $k[\mathcal{C}]=k[x_{0},\ldots,x_{n}]/I_{\mathcal{A}}$, where $\mathcal{S}=\mathcal{S}_{\mathcal{A}}$ denotes the monoid spanned by $\mathcal{A}$. The projective monomial curve $\mathcal{C}$ is said to be arithmetically Cohen-Macaulay (resp. Gorenstein) if the ring $k[\mathcal{C}]$ is Cohen-Macaulay (resp. Gorenstein).
The projective curve $\mathcal{C}$ has two affine charts, $\mathcal{C}_{1}=\{(u^{a_{1}},\ldots,u^{a_{n}})\in\mathbb{A}_{k}^{n}\,|\,u\in k\}$ and $\mathcal{C}_{2}=\{(v^{d-a_{0}},v^{d-a_{1}},\ldots,v^{d-a_{n-1}})\in\mathbb{A}_{k}^{n}\,|\,v\in k\}$, associated to the sequences $a_{1}<\cdots<a_{n}$ and $d-a_{n-1}<\cdots<d-a_{1}<d-a_{0}$, respectively. The second sequence is sometimes called the dual of the first one. Denote by $\mathcal{S}_{1}:=\mathcal{S}_{\mathcal{A}_{1}}$ the numerical semigroup generated by $\mathcal{A}_{1}=\{a_{1},\ldots,a_{n}\}$. The vanishing ideal of $\mathcal{C}_{1}$ is $I_{\mathcal{A}_{1}}\subset k[x_{1},\ldots,x_{n}]$, and hence, its coordinate ring is the one-dimensional ring $k[\mathcal{C}_{1}]\simeq k[x_{1},\ldots,x_{n}]/I_{\mathcal{A}_{1}}$. Moreover, $I_{\mathcal{A}}$ is the homogenization of $I_{\mathcal{A}_{1}}$ with respect to the variable $x_{0}$. Similarly, denoting by $\mathcal{S}_{2}:=\mathcal{S}_{\mathcal{A}_{2}}$ the numerical semigroup generated by $\mathcal{A}_{2}:=\{d-a_{0},d-a_{1},\ldots,d-a_{n-1}\}$, the vanishing ideal of $\mathcal{C}_{2}$ is $I_{\mathcal{A}_{2}}\subset k[x_{0},\ldots,x_{n-1}]$, its coordinate ring is $k[\mathcal{C}_{2}]\simeq k[x_{0},\ldots,x_{n-1}]/I_{\mathcal{A}_{2}}$, and $I_{\mathcal{A}}$ is the homogenization of $I_{\mathcal{A}_{2}}$ with respect to $x_{n}$.
@@
There are several sources of motivation for this work. In [8] the authors completely describe the minimal graded free resolution of affine monomial curves defined by an arithmetic sequence $a_{1}<\cdots<a_{n}$ and, in particular, in [8, Thm. 4.7], they deduce that the Cohen-Macaulay type of these curves, i.e., the last Betti number in the Betti sequence, is given by the only integer $t\in\{1,\ldots,n-1\}$ such that $t\equiv a_{1}-1\ ({\rm mod}\ n-1)$. Interestingly, later in [1, Thm. 2.13], exactly the same result is obtained for projective monomial curves defined by an arithmetic sequence. Moreover, some computer-assisted experiments using [6] showed that in some cases the whole Betti sequences coincide, while in other cases they do not.
@@
    For $4<5<6<7<8$, the Betti sequences of $k[\mathcal{C}_{1}]$ and $k[\mathcal{C}]$ are both $(1,7,14,11,3)$.
@@
    For $4<5<6<7<8<9$, the Betti sequence of $k[\mathcal{C}_{1}]$ is $(1,8,21,25,14,3)$, while the Betti sequence of $k[\mathcal{C}]$ is $(1,12,25,25,14,3)$. Observe that, in this example, the Cohen-Macaulay type of both $k[\mathcal{C}]$ and $k[\mathcal{C}_{1}]$ is 3, as expected by [8, Thm. 4.7] and [1, Thm. 2.13], but the two Betti sequences do not coincide.
@@ 1. Apery sets and their poset structure
Let $d\in\mathbb{Z}^{+}$ and $a_{0}=0<a_{1}<\cdots<a_{n}=d$ be a sequence of relatively prime integers. For each $i=0,\ldots,n$, set $\mathbf{a}_{i}:=(a_{i},d-a_{i})\in\mathbb{N}^{2}$, and consider the three sets $\mathcal{A}_{1}=\{a_{1},\ldots,a_{n}\}$, $\mathcal{A}_{2}=\{d,d-a_{1},\ldots,d-a_{n-1}\}$ and $\mathcal{A}=\{\mathbf{a}_{0},\ldots,\mathbf{a}_{n}\}\subset\mathbb{N}^{2}$. We denote by $\mathcal{C}\subset\mathbb{P}_{k}^{\,n}$ the projective monomial curve defined by $\mathcal{A}$ as defined in the introduction, and by $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ its affine charts, i.e., the affine monomial curves given by $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, respectively. We denote the vanishing ideal of $\mathcal{C}_{i}$ by $I_{\mathcal{A}_{i}}$ for $i=1,2$ and the vanishing ideal of $\mathcal{C}$ by $I_{\mathcal{A}}$. Consider $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ the numerical semigroups generated by $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ respectively, and $\mathcal{S}$ the monoid spanned by $\mathcal{A}$ that we call the homogenization of $\mathcal{S}_{1}$ (with respect to $d$).
@@
As observed before stating the proposition, ()($\Rightarrow$)( ⇒ ) always holds. Let us prove ()($\Leftarrow$)( ⇐ ) when $\mathcal{C}$ is arithmetically Cohen-Macaulay. Since $(y_{1},y_{2}),(z_{1},z_{2})\in{\rm AP}_{\mathcal{S}}$, one has that $y_{1},z_{1}\in{\rm Ap}_{1}$, $y_{2},z_{2}\in{\rm Ap}_{2}$ by Proposition 1.1 c, and $y_{1}+y_{2}\equiv z_{1}+z_{2}\equiv 0\pmod{d}$. Assume that $y_{1}\leq_{1}z_{1}$ and $y_{2}\leq_{2}z_{2}$, then $s_{1}:=z_{1}-y_{1}\in\mathcal{S}_{1}$ and $s_{2}:=z_{2}-y_{2}\in\mathcal{S}_{2}$. Indeed, $s_{i}\in{\rm Ap}_{i}$ for $i=1,2$; otherwise, $z_{i}\notin{\rm Ap}_{i}$. Since $s_{1}+s_{2}=z_{1}+z_{2}-y_{1}-y_{2}\equiv 0\pmod{d}$, then $(s_{1},s_{2})\in\mathcal{S}$ by Proposition 1.1 d, and we are done. ∎
@@ 2. Betti numbers of affine and projective monomial curves
Recall that $I_{\mathcal{A}_{1}}\subset k[x_{1},\ldots,x_{n}]$ and $I_{\mathcal{A}}\subset k[x_{0},\ldots,x_{n}]$ are the vanishing ideals of $\mathcal{C}_{1}$ and $\mathcal{C}$ respectively. When $\mathcal{C}$ is arithmetically Cohen-Macaulay, ${\rm pd}(k[\mathcal{C}])={\rm pd}(k[\mathcal{C}_{1}])$. Moreover, by Proposition 2.2 b, in this case, one has that $|{\rm AP}_{\mathcal{S}}|=|{\rm Ap}_{1}|=d$. The main result in this section is Theorem 2.1 where we give a sufficient condition in terms of the poset structures of the Apery sets ${\rm Ap}_{1}$ and ${\rm AP}_{\mathcal{S}}$ for the Betti sequences of $k[\mathcal{C}_{1}]$ and $k[\mathcal{C}]$ to coincide.
@@
Proof.
We will use Propositions 2.2 and 2.3 below. Proposition 2.2 b provides a characterization of when the two posets $({\rm AP},\leq_{\mathcal{S}})$ and $({\rm Ap}_{1},\leq_{1})$ are isomorphic. This condition is then related in Proposition 2.3 to a property on the reduced Gröbner basis of $I_{\mathcal{A}_{1}}$ with respect to the degree reverse lexicographic order. By applying a recent result of Saha, Sengupta and Srivastava [15, Thm. 4.1], our result follows. ∎
@@ 3. Improving Vu’s bound for equality of Betti numbers of a projective monomial curve and its projection
Take a sequence of nonnegative integers $0=c_{1}<\cdots<c_{n}$, not necessarily relatively prime, and consider, for all $j>0$, the shifted set of integers $\mathcal{A}_{1}^{j}=\{c_{1}+j,\ldots,c_{n}+j\}$, and the semigroup $\mathcal{S}_{1}^{j}$ generated by the sequence $a_{0}:=0<a_{1}:=c_{1}+j<\dots<a_{n}:=c_{n}+j$. Herzog and Srinivasan conjectured that the Betti numbers of $k[\mathcal{S}_{1}^{j}]$ eventually become periodic with period $c_{n}$. In [19], Vu provides a proof of this conjecture together with an explicit value $N$ such that this periodic behavior occurs for all $j>N$. One of the key steps in Vu’s argument is [19, Thm. 5.7] where he proves that, for all $j>N$, the Betti numbers of the affine and projective monomial curves defined by $c_{1}+j<\cdots<c_{n}+j$ coincide. Using our results in Section 2, we provide a smaller value of $N$ such that this occurs.
@@
then the result follows from [15, Thm. 4.1].
@@ 4. Construction of Gorenstein projective monomial curves
A numerical semigroup $\mathcal{S}_{1}$ is symmetric if and only if either $b\in\mathcal{S}_{1}$ or $F(\mathcal{S}_{1})-b\in\mathcal{S}_{1}$ for all $b\in\mathbb{N}$. Kunz proved in [13] that $k[\mathcal{C}_{1}]$ is Gorenstein if and only if $\mathcal{S}_{1}$ is symmetric. In this section we show how to construct an arithmetically Gorenstein projective monomial curve from a symmetric numerical semigroup $\mathcal{T}$. We begin with the following result, which provides a necessary and sufficient condition for $\mathcal{C}$ to be arithmetically Gorenstein and is a consequence of the results in [5].

### v3

@@ Introduction
Now consider an integer $d>0$ and a sequence $a_{0}=0<a_{1}<\cdots<a_{n}=d$ of relatively prime integers, i.e., $\gcd(a_{1},\ldots,a_{n})=1$. Denote by $\mathcal{C}$ the projective monomial curve $\mathcal{C}\subset\mathbb{P}_{k}^{\,n}$ of degree $d$ parametrically defined by $x_{i}=u^{a_{i}}v^{d-a_{i}}$ for all $i\in\{0,\ldots,n\}$, i.e., $\mathcal{C}$ is the Zariski closure of
@@
Taking $\mathcal{A}=\{\mathbf{a}_{0},\ldots,\mathbf{a}_{n}\}\subset\mathbb{N}^{2}$ with $\mathbf{a}_{i}=(a_{i},d-a_{i})$ for all $i=0,\ldots,n$, one has that $I_{\mathcal{A}}$ is the vanishing ideal of $\mathcal{C}$, and the coordinate ring of $\mathcal{C}$ is the two-dimensional ring $k[\mathcal{C}]=k[x_{0},\ldots,x_{n}]/I_{\mathcal{A}}\simeq k[\mathcal{S}_{\mathcal{A}}]$, where $\mathcal{S}_{\mathcal{A}}$ denotes the monoid spanned by $\mathcal{A}$. The projective monomial curve $\mathcal{C}$ is said to be arithmetically Cohen-Macaulay (resp. Gorenstein) if the ring $k[\mathcal{C}]$ is Cohen-Macaulay (resp. Gorenstein).
The projective curve $\mathcal{C}$ has two affine charts, $\mathcal{C}_{1}=\{(u^{a_{1}},\ldots,u^{a_{n}})\in\mathbb{A}_{k}^{n}\,|\,u\in k\}$ and $\mathcal{C}_{2}=\{(v^{d-a_{0}},v^{d-a_{1}},\ldots,v^{d-a_{n-1}})\in\mathbb{A}_{k}^{n}\,|\,v\in k\}$, associated to the sequences $a_{1}<\cdots<a_{n}$ and $d-a_{n-1}<\cdots<d-a_{1}<d-a_{0}$, respectively. The second sequence is sometimes called the dual of the first one. Denote by $\mathcal{S}_{1}:=\mathcal{S}_{\mathcal{A}_{1}}$ the submonoid of $\mathbb{N}$ spanned by $\mathcal{A}_{1}$. Since $a_{1},\ldots,a_{n}$ are relatively prime, then $\mathbb{N}\setminus\mathcal{S}_{1}$ is finite and $\mathcal{S}_{1}$ is called a numerical semigroup. The vanishing ideal of $\mathcal{C}_{1}$ is $I_{\mathcal{A}_{1}}\subset k[x_{1},\ldots,x_{n}]$, and hence, its coordinate ring is the one-dimensional ring $k[\mathcal{C}_{1}]=k[x_{1},\ldots,x_{n}]/I_{\mathcal{A}_{1}}\simeq k[\mathcal{S}_{1}]$. Moreover, $I_{\mathcal{A}}$ is the homogenization of $I_{\mathcal{A}_{1}}$ with respect to the variable $x_{0}$. Similarly, denoting by $\mathcal{S}_{2}:=\mathcal{S}_{\mathcal{A}_{2}}$ the numerical semigroup generated by $\mathcal{A}_{2}:=\{d-a_{0},d-a_{1},\ldots,d-a_{n-1}\}$, the vanishing ideal of $\mathcal{C}_{2}$ is $I_{\mathcal{A}_{2}}\subset k[x_{0},\ldots,x_{n-1}]$, its coordinate ring is $k[\mathcal{C}_{2}]=k[x_{0},\ldots,x_{n-1}]/I_{\mathcal{A}_{2}}\simeq k[\mathcal{S}_{2}]$, and $I_{\mathcal{A}}$ is the homogenization of $I_{\mathcal{A}_{2}}$ with respect to $x_{n}$.
@@
There are several sources of motivation for this work. In [8] the authors completely describe the minimal graded free resolution of affine monomial curves defined by an arithmetic sequence $a_{1}<\cdots<a_{n}$, i.e. $a_{i}=a_{1}+(i-1)e$ for some $e\in\mathbb{Z}^{+}$ such that $\gcd(a_{1},e)=1$. In particular, in [8, Thm. 4.7], they deduce that the Cohen-Macaulay type of these curves, i.e., the last Betti number in the Betti sequence, is given by the only integer $t\in\{1,\ldots,n-1\}$ such that $t\equiv a_{1}-1\ ({\rm mod}\ n-1)$. Interestingly, later in [1, Thm. 2.13], exactly the same result is obtained for projective monomial curves defined by an arithmetic sequence. Moreover, some computer-assisted experiments using [6] showed that in some cases the whole Betti sequences coincide, while in other cases they do not.
@@
    For $4<5<6<7<8$, $\mathcal{C}$ is arithmetically Cohen-Macaulay and the Betti sequences of $k[\mathcal{C}_{1}]$ and $k[\mathcal{C}]$ are both $(1,7,14,11,3)$.
@@
    For $4<5<6<7<8<9$, $\mathcal{C}$ is arithmetically Cohen-Macaulay and the Betti sequence of $k[\mathcal{C}_{1}]$ is $(1,8,21,25,14,3)$, while the Betti sequence of $k[\mathcal{C}]$ is $(1,12,25,25,14,3)$. Observe that, in this example, the Cohen-Macaulay type of both $k[\mathcal{C}]$ and $k[\mathcal{C}_{1}]$ is 3, as expected by [8, Thm. 4.7] and [1, Thm. 2.13], but the two Betti sequences do not coincide.
@@ 1. Apery sets and their poset structure
Let $d\in\mathbb{Z}^{+}$ and $a_{0}=0<a_{1}<\cdots<a_{n}=d$ be a sequence of relatively prime integers. For each $i=0,\ldots,n$, set $\mathbf{a}_{i}:=(a_{i},d-a_{i})\in\mathbb{N}^{2}$, and consider the three sets $\mathcal{A}_{1}=\{a_{1},\ldots,a_{n}\}$, $\mathcal{A}_{2}=\{d,d-a_{1},\ldots,d-a_{n-1}\}$ and $\mathcal{A}=\{\mathbf{a}_{0},\ldots,\mathbf{a}_{n}\}\subset\mathbb{N}^{2}$. We denote by $\mathcal{C}\subset\mathbb{P}_{k}^{\,n}$ the projective monomial curve parametrized by $\mathcal{A}$ as defined in the introduction, and by $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ its affine charts, i.e., the affine monomial curves given by $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, respectively. We denote the vanishing ideal of $\mathcal{C}_{i}$ by $I_{\mathcal{A}_{i}}$ for $i=1,2$ and the vanishing ideal of $\mathcal{C}$ by $I_{\mathcal{A}}$. Consider $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ the numerical semigroups generated by $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ respectively, and $\mathcal{S}$ the monoid spanned by $\mathcal{A}$ that we call the homogenization of $\mathcal{S}_{1}$ (with respect to $d$).
@@
As observed before stating the proposition, ()($\Rightarrow$)( ⇒ ) always holds. Let us prove ()($\Leftarrow$)( ⇐ ) when $\mathcal{C}$ is arithmetically Cohen-Macaulay. Since $(y_{1},y_{2}),(z_{1},z_{2})\in{\rm AP}_{\mathcal{S}}$, one has that $y_{1},z_{1}\in{\rm Ap}_{1}$, $y_{2},z_{2}\in{\rm Ap}_{2}$ by Proposition 1.1 c, and $y_{1}+y_{2}\equiv z_{1}+z_{2}\equiv 0\pmod{d}$. Assume that $y_{1}\leq_{1}z_{1}$ and $y_{2}\leq_{2}z_{2}$, then $s_{1}:=z_{1}-y_{1}\in\mathcal{S}_{1}$ and $s_{2}:=z_{2}-y_{2}\in\mathcal{S}_{2}$. Moreover, $s_{i}\in{\rm Ap}_{i}$ for $i=1,2$; otherwise, $z_{i}\notin{\rm Ap}_{i}$. Since $s_{1}+s_{2}=z_{1}+z_{2}-y_{1}-y_{2}\equiv 0\pmod{d}$, then $(s_{1},s_{2})\in\mathcal{S}$ by Proposition 1.1 d, and we are done. ∎
@@ 2. Betti numbers of affine and projective monomial curves
Recall that $I_{\mathcal{A}_{1}}\subset k[x_{1},\ldots,x_{n}]$ and $I_{\mathcal{A}}\subset k[x_{0},\ldots,x_{n}]$ are the vanishing ideals of $\mathcal{C}_{1}$ and $\mathcal{C}$ respectively. When $\mathcal{C}$ is arithmetically Cohen-Macaulay, ${\rm pd}(k[\mathcal{C}])={\rm pd}(k[\mathcal{C}_{1}])$. Moreover, by Proposition 2.2 b, in this case, one has that $|{\rm AP}_{\mathcal{S}}|=|{\rm Ap}_{1}|=d$. The main result in this section is Theorem 2.1 where we give a sufficient condition in terms of the poset structures of the Apery sets ${\rm Ap}_{1}$ and ${\rm AP}_{\mathcal{S}}$ for the Betti sequences of $k[\mathcal{C}_{1}]$ and $k[\mathcal{C}]$ to coincide. We postpone the proof after Propositions 2.2 and 2.3.
@@
Proof of Theorem 2.1.
By Propositions 2.2 and 2.3, the Apery posets $({\rm AP},\leq_{\mathcal{S}})$ and $({\rm Ap}_{1},\leq_{1})$ are isomorphic if and only if the variable $x_{n}$ appears in every non-homogeneous binomial of $\mathcal{G}_{>}$, the reduced Gröbner basis of $I_{\mathcal{A}_{1}}$ with respect to the degrevlex order with $x_{1}>x_{2}>\dots>x_{n}$. By applying a recent result of Saha, Sengupta and Srivastava [15, Thm. 4.1], our result follows. ∎
@@ 3. Improving Vu’s bound for equality of Betti numbers of a projective monomial curve and its projection
Take a sequence of nonnegative integers $0=c_{1}<\cdots<c_{n}$, not necessarily relatively prime, and consider, for all $j>0$, the shifted set of integers $\mathcal{A}_{1}^{j}=\{c_{1}+j,\ldots,c_{n}+j\}$, and the semigroup $\mathcal{S}_{1}^{j}$ generated by the sequence $a_{0}:=0<a_{1}:=c_{1}+j<\dots<a_{n}:=c_{n}+j$. Herzog and Srinivasan conjectured that the Betti numbers of $k[\mathcal{S}_{1}^{j}]$ eventually become periodic with period $c_{n}$. In [19], Vu provides a proof of this conjecture together with an explicit value $N$ such that this periodic behavior occurs for all $j>N$. One of the key steps in Vu’s argument is [19, Thm. 5.7] where he proves that, for all $j>N$, the Betti numbers of the affine and projective monomial curves defined by $c_{1}+j<\cdots<c_{n}+j$ coincide. We provide a smaller value of $N$ such that this occurs.
@@
then the result follows from [15, Thm. 4.1]. Note that this result is true even if the generators of the semigroup are not relatively prime since the defining ideal does not change when we divide them by a common divisor.
@@ 4. Construction of Gorenstein projective monomial curves
A numerical semigroup $\mathcal{S}_{1}$ is symmetric if and only if either $b\in\mathcal{S}_{1}$ or $F(\mathcal{S}_{1})-b\in\mathcal{S}_{1}$ for all $b\in\mathbb{N}$, where $F(\mathcal{S}_{1})=\max\left(\mathbb{N}\setminus\mathcal{S}_{1}\right)$ is the Frobenius number of $\mathcal{S}_{1}$. Kunz proved in [13] that $k[\mathcal{C}_{1}]$ is Gorenstein if and only if $\mathcal{S}_{1}$ is symmetric. In this section we show how to construct an arithmetically Gorenstein projective monomial curve from a symmetric numerical semigroup $\mathcal{T}$. We begin with the following result, which provides a necessary and sufficient condition for $\mathcal{C}$ to be arithmetically Gorenstein and is a consequence of the results in [5].